# Danoliteracy of Generative, Large Language Models

### v1

@@ Abstract
The language technology moonshot moment of Generative, Large Language Models (GLLMs) was not limited to English: These models brought a surge of technological applications, investments and hype to low-resource languages as well. However, the capabilities of these models in languages such as Danish were until recently difficult to verify beyond qualitative demonstrations due to a lack of applicable evaluation corpora. We present a GLLM benchmark to evaluate Danoliteracy, a measure of Danish language and cultural competency, across eight diverse scenarios such Danish citizenship tests and abstractive social media question answering. This limited-size benchmark is found to produce a robust ranking that correlates to human feedback at $\rho\sim 0.8$ with GPT-4 and Claude Opus models achieving the highest rankings. Analyzing these model results across scenarios, we find one strong underlying factor explaining $95\%$ of scenario performance variance for GLLMs in Danish, suggesting a $g$ factor of model consistency in language adaption.
@@ 1 Introduction
Benchmarks shape technologies. By acting as normative guidelines for technology applications, benchmarks imply directions of research and development that ultimately impact users citeliang2022holistic. GLLMs specifically have emerged as a technology with near-universal impact, including lower-resouce languages such as Danish Olsen (2023). If the challenging task of general GLLM evaluation is not extended to low-resource languages, practitioners start from scratch for each model use case, inhibiting practical adoption or possibly resulting in risky, undertested implementations.
We take up the challenge of creating a GLLM evaluation benchmark for Danish, a North Germanic language spoken by about six million people, primarily in the Nordic country of Denmark. As depicted in Figure 1, our approach is to create a compilation of small-scale, diverse evaluation scenarios combined into one general benchmark to reveal model Danoliteracy. By Danoliteracy we refer to the level of GLLM real-world knowledge (RWK), natural language understanding (NLU), and natural language generation (NLG) in Danish.
This paper presents the resulting Danoliterate Benchmark, describing evaluation methods, datasets and model results. We analyze these results with the goals of validating and exploring evaluation methodology. An important part of this analysis is to investigate the feasibility of such evaluation: Does this small-scale, language-specific approach achieve a significant ranking of GLLMs? Even if a non-spurious leaderboard can be discerned from the result, it is not enough to validate the benchmark which might actually show something orthogonal to Danoliteracy. Thus, as a benchmark validation tool, we additionaly present a user survey, collecting the preferences of Danish speakers when interacting with hidden pairs of GLLMs in an arena A/B test setup.
The availability of a suite of meaningful benchmark results allows us to investigate GLLM behavior: Initially, we explore which specific models are most Danoliterate and how different types of GLLMs compare. Beyond that, we are particularly interested in capability consistency across tasks: If a GLLM performs strongly at one Danish use-case scenario, can we then also expect high capabilities in other Danish scenarios with different domains and objectives?
We hope so. If the answer is no, practitioners are without general results to trust, requring a full model re-evaluation for each downstream use. However, if capability consistency is present, we should be able to find a single underlying axis that correlates with performance across diverse scenarios. Such a general dimension of Danoliteracy can be compared to the $g$ factor of general human intelligence Spearman (1904). If one significant, main factor is found, it implies a level of stability that can help guide the expectations of practitioners across GLLM implementations in varying and even novel Danish tasks.
@@
  • Evidence that GPT-4 and Claude Opus models are currently uniquely capable in Danish, outperforming other closed models which in turn beat open-weights models.
@@ 2.2 Low-resource NLP Evaluation
Other approaches focus on one language exclusively in attempts to evaluate GLLM language performance beyond surface-level lexical or syntactical literacy. Using this method, pracitioners can align scenario domain, cultural content, and real-world facts with the setting of the language, though a lack of relevant data can be problematic Liu et al. (2023).
Specifically in Danish, in the comprehensive ScandEval benchmark, packaging scenarios across eight languages divided into NLU and NLG leaderboards, has implemented evaluation on GLLMs in Danish on eight NLG scenarios with some overlap in dataset sources with this work Nielsen (2023).
@@ 2.3 GLLM $g$ Factor
The idea that GLLM performance is strongly correlated across tasks has been noted previously by for example Ilić who carried out factor analysis on the Open LLM Leaderboard and GLUE Wang et al. (2018) with results similar to ours Ilić (2023).
@@ 3.1 Evaluation
Most of the available datasets allowed primarily for testing discriminatory RWK and NLU of the GLLM by requiring it to select between multiple-choice answers. For these multiple-choice scenarios, frequency of generating the correct option number was reported as model accuracy.
Two metrics were used for NLG. First, summarization was implemented using a similarity score between model summary and a reference summary $\operatorname{s}{\left(\mathcal{T}_{\mathcal{M}},\mathcal{T}_{\text{ref}}\right)}$. Secondly, we implemented abstractive question answering tasks for datasets $D$ where each question had not just one correct answer but a corresponding set of reference, human-generated answers. This was done by scoring GLLMs using the frequency with which generated answers were the odd-one-out, defined by the lowest total similarity to all possible answers $T=\{T_{\mathcal{M}}\}\cup\{\mathcal{T}_{\text{ref, i}}\}_{i=1..k}$ as shown in Eq. 1. For similarity scores $s$, the BERT score algorithm Zhang* et al. (2020) based on the DFM Encoder Large Enevoldsen et al. (2022) was used.
@@
Finally, few-shot named entity recognition (NER) was implemented for GLLMs using 3-shot prompting and the GPT-NER multiple queries idea Wang et al. (2023). Here, word-level entity class predictions were aligned and the standard NER micro average F1 scores were calculated using the SeqEval framework Nakayama (2018).
@@
Figure 3: Model Danoliteracy Index across all scenarios for top performers. Two model nodes are connected iff. the bootstrapping procedure could not reveal significant benchmark performance difference at $\alpha=0.05$. Claude Opus and the GPT 4 family models are consistent winners.
3.2 Datasets
Real-world Knowledge
  1. 1.
  2. 2.
    HyggeSwag is a novel manual translation of 125 HellaSwag Zellers et al. (2019) ActivityNet Caba Heilbron et al. (2015) questions testing commonsense natural language inference as a multiple-choice task to pick the only completion consistent with real-world logic.
  3. 3.
Free NLG
  1. 4
    #twitterhjerne is a novel abtractive question-answering dataset containing 78 anonymized question tweets from the Danish hashtag of that name, translated to Twitter Brain, where users ask the social media hive mind for help, input or recommendations. For each question tweet, 2-9 reference answer tweets were extracted making it possible to use the $\text{score}_{\text{o1o}}$ metric (1).
  2. 5
Figure 4: Selected model normalized results across the eight scenarios divided in three categories as described in Section 3.2. Claude Opus is overtaken by GPT-4 on the NER task but wins on an NLG task. LlaMa 3 70B, the SOTA open-weights model, lacks behind on NLU and knowledge-based tasks. A Danish-specialized model with only 1.1B parameters, DanskGPT-tiny Chat, benchmarks well in NLG but fails on knowledge and understanding.
Figure 5: The non-normalized metric scores across evaluation scenarios for two models that were judged highly according to human feedback. Uncertainties are 95% confidence intervals according to the bootstrapping procedure and the micro-average is displayed for each model.
NLU Tasks
  1. 6
  2. 7
  3. 8
All datasets are released on the Huggingface Datasets Hub with dataset cards [^1] Datasets can be found on danoliterate.compute.dtu.dk/Scenarios except for the two small datasets extracted from CRR which require practitioners to re-run data collection for personal use. More details on dataset licensing and collection as well as as data examples can be found in Appendix B.
@@ 3.3 Models
Both local prediction of open-weights models and API access to externally hosted GLLMs were implemented. 54 autoregressive, decoder language models trained for general text generation were included. Models were tested if we saw any reason to suspect a degree of Danoliteracy, thus including multilingual models with possibly small amounts of Danish training data as well as other mainland Scandinavian monolingual models but excluding strictly English-only models. Both base and instruct-tuned models were evaluated. All model generation was performed with greedy decoding and with a maximum number of generated tokens of 256. The models run locally ranged in sizes from 124M parameters to 13B parameters, resulting in a total project GPU use of $\sim 100$ hours on a single Nvidia H100.
3.4 Human Feedback
For a subset of 18 instruct-tuned models, we have set up a parallel study to collect human judgment on model performance. Volunteers were presented with a anonymized pair of models and were asked to report their preferred model. This was done based on side-by-side model answers on at least three prompts selected by the volunteer from a pool of 100 prompt examples. Prompt selection was chosen independently of the Danoliterate Benchmark by creating one Danish prompt for each of 100 popular generative AI use-cases according to Zao-Sanders Zao-Sanders (2024). The study is ongoing: At the time of writing, 469 responses were analyzed. More details on data collection and analysis can be seen in Appendix C.
@@ 3.5 The Danoliterate Framework
A modular, open-source evaluation framework was implemented in Python, using Huggingface Transformers Wolf et al. (2019) and Datasets Lhoest et al. (2021) as central tools as well as Hydra Yadan (2019) and a Weights and Biases-integration Biewald (2020) for structuring experimentation. This framework, danoliterate, is relased on GitHub222 github.com/sorenmulli/danoliterate under the MIT License.
@@ 4 Results
Figure 7: Model Danish capabilities based on human feedback. Values are normalized Bradley-Terrey coefficients where two models are connected if the coefficiencts are not significantly different in the ranking model.
@@ 4.1 Benchmark Feasibility
To quantify benchmark noise, we implemented blocked bootstrapping, resampling all examples with replacement and aggregating all $N=8$ scenario scores for each of the $M=54$ models. For each $K=10,000$ bootstrap samples, the $M\times N$ model scenario results were aggregated into one overall Danoliteracy Index for each model $d_{\mathcal{M}}$. This index was computed by considering one scenario at a time, assigning to the winner index 100 and to the lowest-scoring model index 0 with a linear scaling between the two. The micro average across the $N$ scenarios is reported as the resulting Danoliteracy Index of the sample.
@@
However, robust separation is visible for some models, providing basis for statements like ”The different GPT-4 models benchmark at the same level but clearly perform better than GPT-3 models” or ”The bigger the Claude 3, the better the performance – but even the cheap Haiku version performs at GPT-3.5 level” or ”Small, Danish-specialized models like Heidrun can perform at LlaMa 2 70B level but LlaMa 3 has moved the SOTA for open-weights models in Danish”. This signal allows us to learn more about reasons for model performance to which we return in the next section.
First, we turn to the important question of validity: We see a robust benchmark signal resulting a significant ranking but must question the meaning of the signal. One superficial indication of a meaningful signal is that, as expected, the ranking correlates significantly with model parameter counts [^4] Here, only 38 open models with known parameter counts were considered. $\alpha=5\%$ confidence interval: $[0.37;0.78]$. $\rho\sim 0.6$. However, more importantly: We find that it does correlate with the preliminary results of our Danish human judgement survey.
@@
Thus, the results from our monolingual scenario compilation approach differ to those from English benchmarks while importantly, showing high correspondence to judgments made by Danish speakers.
@@ 4.2 Model Outcomes
The substantial requirements of dataset and model scale as well as creation of instruct datasets might explain why nationally anchored organisations have not been able to come up with Danish-first models competing with the multilingual behemoths. Such language-general models have the advantage of linguistic and factual knowledge enhancement across training languages which also benefit them in the monolingual setting.
@@ 4.3 Capability Dimensionality
The previous analysis mainly considered the aggregated benchmark results across scenarios. What is going on at the scenario level? While different model capability profiles can be seen, exemplified in Figure 4 and Figure 5, the main first impression is that model performance at one benchmark scenario strongly predicts performance at other scenarios: One principal component explains 77% of the model result variance across the eight scenarios.
@@
Danoliterate $54\times 8$ $95\%$ $5\%$ 1 1
ScandEval Da. $147\times 8$ $91\%$ $9\%$ 1 2
ScandEval Full $147\times 24$ $80\%$ $13\%$ 2 2
HELM Lite $78\times 10$ $76\%$ $21\%$ 2 2
OpenLLM $1602\times 6$ $95\%$ $5\%$ 1 2
Table 1: How much variance did the first and second factors in EFA explain for the Danoliterate Benchmark as well as the ScandEval benchmark, both full and Danish subset Nielsen (2023), English benchmarks HELM Lite Liang et al. (2022) and OpenLLM Beeching et al. (2023) as of October 2024. Leaderboard dimensionality, model count $\times$ scenario count, is presented along with suggested signficant factor count by the Kaiser criterion and Parallel Analysis: All benchmarks have an important first component.
@@ 5 Conclusions
Based on the ability to robustly discover model groupings at different Danish capability levels and correlate these rankings with human feedback, we conclude that a scenario compilation approach can meaningfully reveal GLLM capabilities. We show that, in Danish, open-weights GLLMs currently lack behind large, closed, multilingual, instruct-tuned models, such as GPT-4 and Claude Opus.
For our evaluation setup, we observe one underlying factor in model capability across the diverse test scenarios. This observation is corroborated by similar structures in other Danish, English and multilingual scenario compilations which we consider a positive result for low-resource evaluation: By using curated and language-specific scenarios, the general landscape of GLLM capabilities for a given low-resource language can be meaningfully inferred even if resources limit the scale.

### v2

# Danoliteracy of Generative Large Language Models

@@ Abstract
The language technology moonshot moment of Generative Large Language Models (GLLMs) was not limited to English: These models brought a surge of technological applications, investments, and hype to low-resource languages as well. However, the capabilities of these models in languages such as Danish were, until recently, difficult to verify beyond qualitative demonstrations due to a lack of applicable evaluation corpora. We present a GLLM benchmark to evaluate Danoliteracy, a measure of Danish language and cultural competency across eight diverse scenarios such as Danish citizenship tests and abstractive social media question answering. This limited-size benchmark was found to produce a robust ranking that correlates to human feedback at $\rho\sim 0.8$ with GPT-4 and Claude Opus models achieving the highest rankings. Analyzing these model results across scenarios, we find one strong underlying factor explaining $95\%$ of scenario performance variance for GLLMs in Danish, suggesting a $g$ factor of model consistency in language adaptation.
@@ 1 Introduction
Benchmarks shape technologies. By acting as normative guidelines for technology applications, benchmarks imply directions of research and development that ultimately impact usersLiang et al. (2022). GLLMs specifically have emerged as a technology with near-universal impact, including lower-resource languages such as Danish Olsen (2023). If the challenging task of general GLLM evaluation is not extended to low-resource languages, practitioners start from scratch for each model use case, inhibiting practical adoption or possibly resulting in risky, undertested implementations.
We take up the challenge of creating a GLLM evaluation benchmark for Danish, a North Germanic language spoken by 6 million people, primarily in the Nordic country of Denmark. As depicted in Figure 1, our approach is to create a compilation of small-scale, diverse evaluation scenarios combined into one general benchmark to reveal model Danoliteracy. By Danoliteracy we refer to the level of GLLM real-world knowledge (RWK), natural language understanding (NLU), and natural language generation (NLG) in Danish.
This paper presents the resulting Danoliterate Benchmark, describing evaluation methods, datasets and model results. We analyze these results with the goals of validating and exploring evaluation methodology. An important part of this analysis is to investigate the feasibility of such evaluation: Does this small-scale, language-specific approach achieve a significant ranking of GLLMs? Even if a non-spurious leaderboard can be discerned from the result, it is not enough to validate the benchmark which might actually show something orthogonal to Danoliteracy. Thus, as a benchmark validation tool, we additionally present a user survey, collecting the preferences of Danish speakers when interacting with hidden pairs of GLLMs in an arena A/B test setup.
The availability of a suite of meaningful benchmark results allows us to investigate GLLM behavior: Initially, we explore which specific models are most Danoliterate and how different types of GLLMs compare. Beyond that, we are particularly interested in capability consistency across tasks: If a GLLM performs strongly in one Danish use-case scenario, does this performance generalize to other Danish scenarios across different domains and objectives?
We hope so. If the answer is no, practitioners are without general results to trust, requiring a full model re-evaluation for each downstream use. However, if capability consistency is present, we should be able to find a single underlying axis that correlates with performance across diverse scenarios. Such a general dimension of Danoliteracy can be compared to the $g$ factor of general human intelligence Spearman (1904). If one significant, main factor is found, it implies a level of stability that can help guide the expectations of practitioners across GLLM implementations in varying and even novel Danish tasks.
@@
  • Evidence that GPT-4 and Claude Opus models are currently uniquely capable in Danish, outperforming other closed models which in turn overcome open-weights models.
@@ 2.2 Low-resource NLP Evaluation
Other approaches focus on one language exclusively in attempts to evaluate GLLM language performance beyond surface-level lexical or syntactical literacy. Using this method, practitioners can align scenario domain, cultural content, and real-world facts with the setting of the language, though a lack of relevant data can be problematic Liu et al. (2023).
Specifically in Danish, the comprehensive ScandEval benchmark, which packages scenarios across eight languages divided into NLU and NLG leaderboards, implements evaluation on GLLMs in Danish on eight NLG scenarios with some overlap in dataset sources with this work Nielsen (2023).
@@ 2.3 GLLM $g$ Factor
The idea that GLLM performance is strongly correlated across tasks has been noted previously by for example Ilić who carried out factor analysis on the Open LLM Leaderboard and GLUE, Wang et al. (2018) obtaining results similar to ours Ilić (2023).
@@ 3 Methods
3.1 Datasets
Real-world Knowledge
  1. 1.
  2. 2.
    HyggeSwag is a novel manual translation [^1] The text was translated by the authors with each translation being validated by another author completing the inference task. of 125 HellaSwag Zellers et al. (2019) ActivityNet Caba Heilbron et al. (2015) questions testing commonsense natural language inference as a multiple-choice task to pick the only completion consistent with real-world logic.
  3. 3.
Free NLG
  1. 4
    #twitterhjerne is a novel abstractive question-answering dataset containing 78 anonymized question tweets from the Danish hashtag of that name, translated to Twitter Brain, where users ask the social media hive mind for help, input or recommendations. For each question tweet, 2-9 reference answer tweets were extracted making it possible to use the $\text{score}_{\text{o1o}}$ metric (1).
  2. 5
Figure 2: Selected model normalized results across the eight scenarios divided into three categories as described in Section 3.1. Claude Opus is overtaken by GPT-4 on the NER task but wins on an NLG task. LlaMa 3 70B, the SOTA open-weights model, lags behind on NLU and knowledge-based tasks. A Danish-specialized model with only 1.1B parameters, DanskGPT-tiny Chat, benchmarks well in NLG but fails on knowledge and understanding.
Figure 3: The non-normalized metric scores across evaluation scenarios for two models that were judged highly according to human feedback. Uncertainties are 95% confidence intervals according to the bootstrapping procedure and the micro-average is displayed for each model. k
NLU Tasks
  1. 6
  2. 7
  3. 8
All datasets are released on the Huggingface Datasets Hub with dataset cards [^2] Datasets can be found on danoliterate.compute.dtu.dk/Scenarios except for the two small datasets extracted from CRR which require practitioners to re-run data collection for personal use. More details on dataset licensing and collection as well as data examples can be found in Appendix B.
@@ 3.2 Evaluation
Most of the available datasets allowed primarily for testing discriminative RWK and NLU of the GLLM by requiring it to select between multiple-choice answers. For these multiple-choice scenarios, frequency of generating the correct option number was reported as model accuracy.
Two metrics were used for NLG. First, summarization was implemented using a similarity score between model summary and a reference summary $\operatorname{s}{\left(\mathcal{T}_{\mathcal{M}},\mathcal{T}_{\text{ref}}\right)}$. Secondly, we implemented abstractive question answering tasks for the specific type of dataset $D$ where each question has not just one correct answer but a corresponding set of reference, human-generated answers. This was done by scoring GLLMs using the frequency with which generated answers were the odd-one-out, defined by the lowest total similarity to all possible answers $T=\{T_{\mathcal{M}}\}\cup\{\mathcal{T}_{\text{ref, i}}\}_{i=1..k}$ as shown in Eq. 1. For similarity scores $s$, the BERT score algorithm Zhang* et al. (2020) based on the DFM Encoder Large Enevoldsen et al. (2022) was used.
@@
Finally, few-shot named entity recognition (NER) was implemented for GLLMs using 3-shot prompting and the GPT-NER multiple queries idea Wang et al. (2023). Here, word-level entity class predictions were aligned and the standard NER micro-average F1 scores were calculated using the SeqEval framework Nakayama (2018).
@@
Figure 5: Model Danoliteracy Index across all scenarios for top performers. Two model nodes are connected iff the bootstrapping procedure could not reveal significant benchmark performance difference at $\alpha=0.05$. Together with the special o1 model, Claude Opus and the GPT 4 family models are consistent winners.
@@ 3.3 Models
Both local prediction of open-weights models and API access to externally hosted GLLMs were implemented. 54 autoregressive, decoder language models trained for general text generation were included. Models were tested if we saw any reason to suspect a degree of Danoliteracy, thus including multilingual models with possibly small amounts of Danish training data as well as other Mainland Scandinavian monolingual models but excluding strictly English-only models. Both base and instruct-tuned models were evaluated. All model generation was performed with greedy decoding and with a maximum number of generated model tokens of 256. OpenAI’s o1 model was allowed to generate internal tokens freely. The models run locally ranged in sizes from 124M parameters to 13B parameters, resulting in a total project GPU use of $\sim 100$ hours on a single Nvidia H100.
@@ 3.4 The Danoliterate Framework
A modular, open-source evaluation framework was implemented in Python, using Huggingface Transformers Wolf et al. (2019) and Datasets Lhoest et al. (2021) as central tools as well as Hydra Yadan (2019) and a Weights and Biases-integration Biewald (2020) for structuring experimentation. This framework, danoliterate, is released on GitHub333 github.com/sorenmulli/danoliterate under the MIT License.
@@
3.5 Human Feedback
For a subset of 18 instruct-tuned models, we have set up a parallel study to collect human judgment on model performance. Volunteers were presented with a anonymized pair of models and were asked to report their preferred model. This was done based on side-by-side model answers on at least three prompts selected by the volunteer from a pool of 100 prompt examples. Prompt selection was chosen independently of the Danoliterate Benchmark by creating one Danish prompt for each of 100 popular generative AI use-cases according to Zao-Sanders Zao-Sanders (2024). The study is ongoing: At the time of writing, 477 responses were analyzed. More details on data collection and analysis can be seen in Appendix C.
@@ 4 Results
Figure 7: Model Danish capabilities based on human feedback. Values are normalized Bradley-Terrey coefficients where two models are connected if the coefficients are not significantly different in the ranking model described in Appendix C.3.
@@ 4.1 Benchmark Feasibility
To quantify benchmark noise, we implemented blocked bootstrapping, resampling all examples with replacement and aggregating all $N=8$ scenario scores for each of the $M=83$ models. For each $K=10,000$ bootstrap samples, the $M\times N$ model scenario results were aggregated into one overall Danoliteracy Index for each model $d_{\mathcal{M}}$. This index was computed by considering one scenario at a time, assigning to the winner index 100 and to the lowest-scoring model index 0 with a linear scaling between the two. The micro-average across the $N$ scenarios is reported as the resulting Danoliteracy Index of the sample.
@@
However, robust separation is visible for some models, providing basis for statements like ”The different GPT-4 models benchmark at the same level but clearly perform better than GPT-3 models” or ”The bigger the Claude 3, the better the performance – but even the cheap Haiku version performs at GPT-3.5 level” or ”Small, Danish-specialized models like Heidrun can perform at LlaMa 2 70B level but LlaMa 3 has moved the SOTA for open-weights models in Danish”. This signal allows us to learn more about reasons for model performance which we explore in the next section.
First, we turn to the important question of validity: We see a robust benchmark signal resulting in a significant ranking but must question the meaning of the signal. One superficial indication of a meaningful signal is that, as expected, the ranking correlates significantly with model parameter counts [^5] Here, only 38 open models with known parameter counts were considered. $\alpha=5\%$ confidence interval: $[0.37;0.78]$. $\rho\sim 0.6$. However, more importantly: We find that it does correlate with the preliminary results of our Danish human judgement survey.
@@
Thus, the results from our monolingual scenario compilation approach differ from those from English benchmarks while importantly, showing high correspondence to judgments made by Danish speakers.
@@ 4.2 Model Outcomes
The substantial requirements of dataset and model scale as well as creation of instruct datasets might explain why nationally anchored organizations have not been able to come up with Danish-first models competing with the multilingual behemoths. Such multilingual models have the advantage of linguistic and factual knowledge enhancement across training languages which also benefit them in the monolingual setting.
@@ 4.3 Capability Dimensionality
The previous analysis primarily considered the aggregated benchmark results across scenarios. What is going on at the scenario level? While different model capability profiles can be seen, exemplified in Figure 2 and Figure 3, the main first impression is that model performance at one benchmark scenario strongly predicts performance at other scenarios: One principal component explains 75% of the model result variance across the eight scenarios.
@@
Danoliterate $83\times 8$ $93\%$ $7\%$ 1 1
ScandEval Da. $199\times 8$ $88\%$ $12\%$ 1 2
ScandEval Full $199\times 24$ $77\%$ $16\%$ 2 2
HELM Lite $90\times 10$ $78\%$ $19\%$ 2 2
OpenLLM $2859\times 6$ $97\%$ $3\%$ 1 2
Table 1: How much variance did the first and second factors in EFA explain for the Danoliterate Benchmark as well as the ScandEval benchmark, both full and Danish subset Nielsen (2023), English benchmarks HELM Lite Liang et al. (2022) and OpenLLM Beeching et al. (2023) as of January 2025. Leaderboard dimensionality, model count $\times$ scenario count, is presented along with suggested significant factor count by the Kaiser criterion and Parallel Analysis: All benchmarks have an important first component.
@@ 5 Conclusions
Based on the ability to robustly discover model groupings at different Danish capability levels and correlate these rankings with human feedback, we conclude that a scenario compilation approach can meaningfully reveal GLLM capabilities. We show that, in Danish, open-weights GLLMs currently lag behind large, closed, multilingual, instruct-tuned models, such as GPT-4 and Claude Opus.
For our evaluation setup, we observe one underlying factor in model capability across the diverse test scenarios. This observation is supported by similar structures in other Danish, English and multilingual scenario compilations which we consider a positive result for low-resource evaluation: By using curated and language-specific scenarios, the general landscape of GLLM capabilities for a given low-resource language can be meaningfully inferred even if resources limit the scale.